# Higher-Order Staircase Codes

### v1

@@ Abstract
We generalize staircase codes and tiled diagonal zipper codes, preserving their key properties while allowing each coded symbol to be protected by arbitrarily many component codewords rather than only two. This generalization arises from the marriage of two distinct combinatorial objects: difference triangle sets and finite-geometric nets, which have typically been applied separately to code design. We demonstrate one realization of the proposed codes, obtaining powerful, high-rate, low-error-floor, and low-complexity coding schemes based on simple iterative syndrome-domain decoding of coupled Hamming component codes. We anticipate that the proposed codes improve performance–complexity–latency tradeoffs in high-throughput communications applications, most notably fiber-optic, in which classical staircase codes and zipper codes have been applied.
@@ I Overview
Staircase codes [2] and related product-like codes [3] have provided a highly-competitive paradigm for energy-efficient, high-rate, high-throughput, and high-reliability, error control coding. Such codes use syndrome-domain iterative algebraic or bounded-distance decoding of algebraic component codes to enable high decoder throughputs, while maintaining relatively low internal decoder data-flow. Furthermore, their block-convolutional structures translate to natural pipelining and parallelism, the combination of which is essential for efficient hardware designs. Moreover, their typically combinatorial constructions ensure control over error floors which must usually be below $10^{-15}$ in the relevant applications. Most notably, such codes are used in fiber-optic communication [4, Ch. 7], but are also relevant to flash storage applications [5] which increasingly demand similar code characteristics.
@@
In this paper, we provide a fruitful generalization of staircase codes which arises from the combination of difference triangle sets (DTSs) [6, Part VI: Ch. 19] and finite-geometric nets [6, Part III: Ch. 3]. Both DTSs and nets are well-studied combinatorial objects found in the standard reference [6] and both have been separately applied to code design, e.g., in [7, 8] and [9, 10] respectively. In these examples, DTSs lead naturally to convolutional codes while nets lead naturally to block codes. It is unsurprising then that their combination leads to a generalization of staircase codes which themselves combine aspects of block and convolutional coding.
@@
We will proceed to summarize, for the reader’s convenience, the key construction and observations of this paper. A higher-order staircase code is determined by three intertwined objects: a DTS, a net, and a component code. Each of these objects is characterized by two parameters, the meanings of which will be defined later. The relationships between these parameters and properties of the resulting code are as follows: Given an $(L,M)$-DTS, an $(M+1,S/L)$-net, and a component code of length $(M+1)S$ and dimension $(M+1)S-r$, one obtains a rate $1-r/S$ spatially-coupled code in which every symbol is protected by $M+1$ component codewords and any pair of distinct component codewords share at most one symbol. Alternatively, the Tanner graph of the code is an infinite semiregular $4$-cycle-free bipartite graph with variable node degree $M+1$ and generalized constraint node degree $(M+1)S$. The relevance of the parameter $L$ is that increasing it can reduce the encoding and decoding memory at the cost of reduced parallelism while keeping all of the aforementioned code properties completely unchanged.
@@
The proposed code family recovers as particular cases some existing codes and interpolates between them:
  • When $L=M=1$, the classical staircase codes of [2] are recovered.
@@
  • When $L>1$ and $M=1$, the tiled diagonal zipper codes of [3] are recovered.
@@
  • When $S=L$ and $r=1$, we recover a recursive (rather than feedforward) version of the self-orthogonal convolutional codes of [7].
@@
When $M>1$ and $L=1$, we obtain a generalization of staircase codes with higher variable degree which was included in a conference version of this work [1]. A similar construction is included in the recently issued US patent [11], which recognizes the role of DTSs in such a generalization of staircase codes but makes no explicit use of finite-geometric nets. Importantly, when $L>1$ and $M>1$, our construction becomes a higher variable degree generalization of tiled diagonal zipper codes [3] which reduces memory and provides flexibility in hardware implementations.
@@
The proposed codes, like their predecessors in [2, 3], combine aspects of convolutional codes, low-density parity-check (LDPC) codes, and algebraic codes. They can be described as an instance of zipper codes[3] or spatially-coupled generalized LDPC (SC-GLDPC) codes [12]. As a result, they admit a multitude of equivalent descriptions the most natural of which will depend strongly on the background of the reader. Without discounting the potential significance of these connections, we will limit our goal in this paper to providing as simple and self-contained a description of the proposed codes as possible.
The remainder of this paper is organized as follows. We begin in Section II with a special case of our construction corresponding to $L=1$ along with key tools and definitions. In Section III, we apply the resulting construction to obtaining powerful, low-error-floor Hamming-based product-like codes. Finally, we provide the complete construction corresponding to $L\geq 1$ in Section IV, before concluding in Section V with suggestions for future work.
II Generalized Staircase Codes with Higher Bit Degree ($L=1$, $M\geq 1$)
@@
In this section, we provide a generalization of staircase codes [2] in which each coded bit (or nonbinary symbol) is protected by arbitrarily many component codewords rather than only two. The goal is to do this while preserving the property of classical staircase codes, essential to error floor performance, that distinct component codewords intersect in at most one bit. This construction will arise from the combination of a $(1,M)$-DTS and an $(M+1,S)$-net. A $(1,M)$-DTS is equivalent to a Golomb ruler of order $M+1$, to be defined shortly, with the definition of a general $(L,M)$-DTS deferred to Section IV where it is needed.
@@ II-A Code Structure
As with classical staircase codes [2], a codeword is an infinite sequence $B_{0},B_{1},B_{2},\dots$ of $S\times S$ matrices satisfying certain constraints. The parameter $S$ is referred to as the sidelength. Matrix entries are from the binary field $\mathbb{F}_{2}$ for concreteness but nonbinary alphabets are straightforwardly accommodated. We assume zero-based indexing of matrices with the $(i,j)$th entry of a matrix $B$ denoted by $(B)_{(i,j)}$ and the index set $\{0,1,\dots,S-1\}$ denoted by $[S]$.
@@
indexed by $k\in[M+1]$ where $M$ is termed the memory parameter. In other words, for each ${k\in[M+1]}$, we have an invertible function $\pi_{k}$ which maps any given $(i,j)\in[S]\times[S]$ to $\pi_{k}(i,j)\in[S]\times[S]$. We further define
@@
for all $(i,j)\in[S]\times[S]$.
Next, we specify $M+1$ distinct integers
$d_{0}<d_{1}<\dots<d_{M}$
which we refer to as a ruler of order $M+1$. Finally, we specify a linear, systematic $t$-error-correcting component code $\mathcal{C}\subseteq\mathbb{F}_{2}^{(M+1)S}$ of length $(M+1)S$ and dimension $(M+1)S-r$ thus having $r$ parity bits. We can without loss of generality assume that $d_{0}=0$ and that $\Pi_{0}(B)=B$ or, equivalently, $\pi_{0}(i,j)=(i,j)$.
@@
which satisfies the constraint that the rows of (1) belong to $\mathcal{C}$.
Therefore, each transmitted block $B_{n}$ contains $S-r$ information columns and $r$ parity columns yielding a rate of
@@
It remains to specify precisely what the permutations and ruler should be. Before proceeding, we provide a graph perspective which formalizes this question. This can be skimmed over by a reader who is solely interested in a description of the proposed codes without their mathematical justification.
@@ II-B Graph Perspective and Scattering
For mathematical convenience, especially in the later parts of this paper, we extend staircase codewords to be bi-infinite sequences
@@
so that the constraint on the rows of (1) in a generalized staircase code can be presumed to hold for all integers $n\in\mathbb{Z}$ rather than only $n\geq d_{M}$. Practically, nothing changes since the all-zero initialization blocks are not transmitted and only a finite number of past blocks are ever needed for encoding.
@@
i.e., that pairs of distinct constraints share at most one variable.
@@ II-C Intra-Block Scattering Permutations via Finite-Geometric Nets
As will be seen shortly, intra-block scattering families of permutations are most easily constructed when $\mathcal{R}$ has sufficiently many elements with invertible differences. Finite fields are an obvious choice since any pair of distinct elements has an invertible difference since all nonzero elements are invertible. We can then take $\mathcal{R}=\mathbb{F}_{q}$, the finite field of order $q$, with $S=q$ a prime power and associate $[S]$ with $\{0,1,\alpha,\alpha^{2},\dots,\alpha^{S-2}\}$ where $\alpha$ is a primitive element of $\mathbb{F}_{q}$. However, a convenient alternative option is to take $\mathcal{R}=\mathbb{Z}_{S}$, the ring of integers modulo $S$, where $S$ is chosen to have a sufficiently large least prime factor denoted $\mathsf{lpf}(S)$. This is because $\{0,1,\dots,\mathsf{lpf}(S)-1\}$ have invertible differences modulo $S$. In particular, we have the following remark:
@@
We now proceed to define $(M+1,S)$-nets which are are well-studied objects in combinatorics and finite geometry with close connections or equivalences to orthogonal arrays, mutually orthogonal Latin squares, transversal designs, and projective planes as discussed in [6].
@@
A large number of constructions and open problems are associated with $(M+1,S)$-nets and can be found in [6]. Fortunately, for our practically-oriented purposes, linear-algebraic permutations over conveniently chosen rings already provide enough flexibility and are straightforward to characterize. In particular, we have the following:
@@
We emphasize that in Examples 1 and 2, we must choose $S$ so that $M\leq\mathsf{lpf}(S)$ otherwise an $(M+1,S)$-net is not formed and the error floor performance suffers. If $S=p$ a prime number, then this condition becomes that $M\leq S$. If $M=S$ is required, then Example 3 provides more flexibility in the admissible values of $S$ which can be a prime power.
@@ II-E Inter-Block Scattering via Golomb Rulers
A set of $M+1$ distinct integers $d_{0}<d_{1}<\dots<d_{M}$, referred to as marks, form a Golomb ruler of length $d_{M}-d_{0}$ and order $M+1$ if and only if all positive differences between marks $d_{k_{2}}-d_{k_{1}}$ with $k_{2}>k_{1}$ (or distances between distinct marks) are distinct.
We again assume that $d_{0}=0$ without loss of generality since translating a ruler by a constant does not alter its distances. Evidently, for a Golomb ruler, the set of distances or positive differences is the set of positive integers
@@
with cardinality $\binom{M+1}{2}=\frac{(M+1)M}{2}$ implying that this is a lower bound on its length $d_{M}$ which is its largest distance. In particular, we have
@@
for any Golomb ruler of order $M+1$ with equality if and only if its set of distances is precisely $\{1,2,\dots,(M+1)M/2\}$.
@@
We provide for the reader’s convenience optimal Golomb rulers of order up to $M+1=10$ in Table I. Observe that the first four optimal Golomb rulers are verified to be optimal by seeing that they meet the lower bound (12). Beyond these cases, optimal Golomb rulers are typically found by computer search. Optimal Golomb rulers of order up to $M+1=20$ are compiled in [6, Part VI: Ch. 19]. While it is an open problem to exhibit optimal Golomb rulers of every order, Golomb rulers of length $d_{M}<(M+1)^{2}$ have been found for all $M+1<65000$ in [14]. Moreover, an explicit construction of Golomb rulers with cubic length in arbitrary order $M+1$ can also be found in [14].

### v5

@@ Abstract
We generalize staircase codes and tiled diagonal zipper codes, preserving their key properties while allowing each coded symbol to be protected by arbitrarily many component codewords rather than only two. This generalization which we term “higher-order staircase codes” arises from the marriage of two distinct combinatorial objects: difference triangle sets and finite-geometric nets, which have typically been applied separately to code design. We demonstrate one possible realization of these codes, obtaining powerful, high-rate, low-error-floor, and low-complexity coding schemes based on simple iterative syndrome-domain decoding of coupled Hamming component codes. We anticipate that the proposed codes could improve performance–complexity–latency tradeoffs in high-throughput communications applications, most notably fiber-optic, in which classical staircase codes and zipper codes have been applied. We consider the construction of difference triangle sets having minimum scope and sum-of-lengths, which lead to memory-optimal realizations of higher-order staircase codes. These results also enable memory reductions for early families of convolutional codes constructed from difference triangle sets.
@@ I Overview
Staircase codes [2] and related product-like codes [3] have provided a highly-competitive paradigm for energy-efficient, high-rate, high-throughput, and high-reliability, error control coding. Such codes use syndrome-domain iterative algebraic or bounded distance decoding of algebraic component codes to enable high decoder throughputs, while maintaining relatively low internal decoder data-flow. Furthermore, their block-convolutional structures translate to natural pipelining and parallelism, the combination of which is essential for efficient hardware designs. Moreover, their typically combinatorial constructions ensure control over error floors which must usually be below $10^{-15}$ in the relevant applications. Most notably, such codes are used in fiber-optic communications [4, Ch. 7], but are also relevant to flash storage applications [5] which increasingly demand similar code characteristics.
@@
In this paper, we provide a fruitful generalization of staircase codes which arises from the combination of difference triangle sets (DTSs) [6, Part VI: Ch. 19] and finite-geometric nets [6, Part III: Ch. 3]. Both DTSs and nets are well-studied combinatorial objects and both have been separately applied to code design, e.g., in [7, 8] and [9, 10] respectively. In these examples, DTSs lead naturally to convolutional codes while nets lead naturally to block codes. It is unsurprising then that their combination leads to a generalization of staircase codes which themselves combine aspects of block and convolutional coding.
@@
We will proceed to summarize at a high level our key construction. A higher-order staircase code is determined by three intertwined objects: a DTS, a net, and a component code. Each of these objects is characterized by two parameters, the meanings of which will be defined later. The relationships between these parameters and properties of the resulting code are as follows: Given an $(L,M)$-DTS, an $(M+1,S/L)$-net, and a component code of length $(M+1)S$ and dimension $(M+1)S-r$, one obtains a rate $1-r/S$ spatially-coupled code in which every symbol is protected by $M+1$ component codewords and any pair of distinct component codewords share at most one symbol. Alternatively, the Tanner graph of the code is an infinite semiregular $4$-cycle-free bipartite graph with variable node degree $M+1$ and generalized constraint node degree $(M+1)S$. The parameter $L$ controls the balance of memory depth against parallelism while keeping all of the aforementioned code properties completely unchanged. A fourth parameter $C$ controls the parallelism obtained by circularly coupling multiple copies of a higher-order staircase code together. The parameters $M$ and $C$ provide mechanisms for improving error floor and threshold performance respectively without altering the component code.
@@
The proposed code family recovers as particular cases some existing code families and interpolates between them, thus providing a unified generalization:
  • When $L=M=C=1$, the classical staircase codes of Smith et al. [2] are recovered.
  • When $L\geq 1$ and $M=C=1$, the tiled diagonal zipper codes of Sukmadji et al. [11] are recovered.
  • When $L\geq 1$, $M=1$, and $C=2$, the OFEC code family [12] is recovered.
@@
  • When $L\geq 1$, $M=1$, and $C\geq 2$, the multiply-chained tiled diagonal zipper codes of Zhao et al. [13] are recovered.
  • When $S/L=M=C=1$, the continuously interleaved codes of Coe [14] are recovered.
@@
  • When $S/L=C=r=1$, a recursive (rather than feedforward) version of the self-orthogonal convolutional codes of Robinson and Bernstein [7] is recovered.
@@
When $M>1$ and $L=1$, we obtain a generalization of staircase codes with higher variable degree which was included in a conference version of this work [1]. A similar construction is included in the recently-issued US patent [15], which recognizes the role of $(1,M)$-DTSs in such a generalization of staircase codes, but makes no explicit use of finite-geometric nets or general $(L,M)$-DTSs. Importantly, when $L>1$ and $M>1$, our construction becomes a higher-variable-degree generalization of tiled diagonal zipper codes [11] which reduces encoding and decoding memory while also providing flexibility in aspects of hardware implementation.
Constructing memory-optimal higher-order staircase codes reduces to a natural extension of the classically studied minimum scope DTS problem [6, Part VI: Ch. 19] in which we additionally seek to minimize a new objective: sum-of-lengths. We study this problem, exhibiting new DTSs that also have implications for earlier work on classical convolutional codes constructed from DTSs.
@@
The proposed codes, like their predecessors in [2, 3], combine aspects of convolutional codes, low-density parity-check (LDPC) codes, and algebraic codes. They can be described as an instance of zipper codes [3], spatially-coupled generalized LDPC (SC-GLDPC) codes [16], or even simply as (highly complex) convolutional codes. However, the proposed codes are very highly structured relative to what is typically admitted under these frameworks. This means that the simplest possible description will be one which is self-contained and does not explicitly appeal to them. Without discounting the potential significance of these connections, we will limit our goal in this paper to providing such a description.
The key advantages of the code structure considered is that it simultaneously enables a large degree of parallelism due to a block-based construction, high performance due to spatial coupling, and control over error floors due to the combinatorial construction. The recovery of established high-throughput coding techniques [2, 11, 12, 14, 13] as special cases is then unsurprising since these properties are precisely what is demanded by such applications.
The remainder of this paper is organized as follows: We begin in Section II with a special case of our construction corresponding to $L=1$ along with key tools and definitions. In Section III, we illustrate one possible realization of the construction of Section II, discussing various practical aspects and a high-throughput simulation technique. Section III also proposes techniques for efficiently decoding systematic Hamming codes without lookup tables (LUTs). Following this, we provide the general, complete construction of higher-order staircase codes corresponding to $L\geq 1$ in Section IV. We follow this with some simulation-based demonstrations in Section V. In Section VI, we study the problem of constructing DTSs that correspond to memory-optimal instances of higher-order staircase codes. A variety of such DTSs are exhibited in the Appendix. Finally, we describe our DTS search technique in Section VII before concluding in Section VIII with suggestions for future work.
II Generalized Staircase Codes with Arbitrary Bit Degree ($L=1$, $M\geq 1$)
@@
In this section, we provide a generalization of staircase codes [2] in which each coded bit (or nonbinary symbol) is protected by arbitrarily many component codewords rather than only two. The goal is to do this while preserving the property of classical staircase codes, essential to error floor performance, that distinct component codewords intersect in at most one bit. This construction will arise from the combination of a $(1,M)$-DTS and an $(M+1,S)$-net. A $(1,M)$-DTS is equivalent to a Golomb ruler of order $M+1$, to be defined shortly, with the definition of a general $(L,M)$-DTS deferred to Section IV where it is needed. Classical staircase codes [2] will be defined as a special case of generalized staircase codes in Section II-F.
@@ II-A Code Structure
As with classical staircase codes [2], a codeword is an infinite sequence $B_{0},B_{1},B_{2},\dots$ of $S\times S$ matrices called blocks satisfying certain constraints. The parameter $S$ is referred to as the sidelength. Matrix entries are from the binary field $\mathbb{F}_{2}$ for concreteness but nonbinary alphabets are straightforwardly accommodated. We assume zero-based indexing of matrices with the $(i,j)$th entry of a matrix $B$ denoted by $(B)_{(i,j)}$ and the index set $\{0,1,\dots,S-1\}$ denoted by $[S]$.
@@
indexed by $k\in[M+1]$ where $M$ is termed the degree parameter. In other words, for each ${k\in[M+1]}$, we have an invertible function $\pi_{k}$ which maps any given $(i,j)\in[S]\times[S]$ to $\pi_{k}(i,j)\in[S]\times[S]$. We further define
@@
for all $(i,j)\in[S]\times[S]$. We will assume that $\Pi_{0}(B)=B$ or, equivalently, that $\pi_{0}(i,j)=(i,j)$, unless otherwise stated. This is typically without loss of generality.
Next, we define rulers:
Definition 1 (Ruler).
A ruler of order $M+1$ is an ordered set of $M+1$ distinct integers $(d_{0},d_{1},\dots,d_{M})$ referred to as marks. A ruler is normalized if $0=d_{0}<d_{1}<\dots<d_{M}$.
We will assume that all rulers are normalized unless otherwise stated. Normalization is achieved by sorting followed by subtraction of the least element from each mark.
Finally, we specify a linear, systematic $t$-error-correcting component code $\mathcal{C}\subseteq\mathbb{F}_{2}^{(M+1)S}$ of length $(M+1)S$ and dimension $(M+1)S-r$ thus having $r$ parity bits. This allows us to provide the following definition for a weak generalized staircase code. The qualifier weak will be dropped once we impose certain conditions on the ruler and permutations.
@@
which satisfies the constraint that the rows of (1) belong to $\mathcal{C}$. Therefore, each transmitted block $B_{n}$ contains $S-r$ information columns and $r$ parity columns yielding a rate of
@@
Figure 1: Visualization of a (classical) staircase code [2] which corresponds to $M=1$, $d_{1}=1$, and $\Pi_{1}(B)=B^{\mathsf{T}}$ (the matrix transpose); rows belong to $\mathcal{C}$.
Figure 2: Visualization of a (weak) generalized staircase code which corresponds to $M=2$, $(d_{1},d_{2})=(1,3)$, $\Pi_{1}(B)=B^{\mathsf{T}}$ (the matrix transpose), and $\Pi_{2}(B)=B^{\pi}$ (some permutation); rows belong to $\mathcal{C}$.
We provide a visualization of the resulting codes in Fig. 1 and Fig. 2 for the cases of $M=1$ and $M=2$ respectively in the style of zipper codes [3]. In these pictures, the rightmost column represents the blocks that are actually transmitted while the columns to its left are delayed and permuted copies that are used to constrain the transmitted blocks.
It remains to specify precisely what the permutations and ruler should be. Before proceeding, we provide a graph perspective which formalizes this question. This can be skimmed on a first reading.
@@ II-B Graph Perspective and Scattering
For mathematical convenience, especially in the later parts of this paper, we extend codewords to be bi-infinite sequences
@@
so that the constraint on the rows of (1) in a weak generalized staircase code can be presumed to hold for all integers $n\in\mathbb{Z}$ rather than only $n\geq d_{M}$. Practically, nothing changes since the all-zero initialization blocks are not transmitted and only a finite number of past blocks are ever needed for encoding.
@@
i.e., that pairs of distinct (generalized) constraints share at most one variable.
@@
Note that this bound is usually—but not always—quite loose and does not address the multiplicity of uncorrectable error patterns. Its purpose is solely to provide a theoretical guarantee for asymptotic control over the error floor rather than to estimate the error floor or prescribe parameters.
@@ II-C Intra-Block Scattering Permutations via Finite-Geometric Nets
As will be seen shortly, intra-block scattering families of permutations are most easily constructed when $\mathcal{R}$ has sufficiently many elements with invertible differences. Finite fields are an obvious choice since any pair of distinct elements has an invertible difference. We can then take $\mathcal{R}=\mathbb{F}_{q}$, the finite field of order $q$, with $S=q$ a prime power and associate $[S]$ with $\{0,1,\alpha,\alpha^{2},\dots,\alpha^{S-2}\}$ where $\alpha$ is a primitive element of $\mathbb{F}_{q}$. However, a convenient alternative option is to take $\mathcal{R}=\mathbb{Z}_{S}$, the ring of integers modulo $S$, where $S$ is chosen to have a sufficiently large least prime factor denoted $\mathsf{lpf}(S)$. This is because $\{0,1,\dots,\mathsf{lpf}(S)-1\}$ have invertible differences modulo $S$. In particular, we have the following remark:
@@
We now proceed to define $(M+1,S)$-nets which are well-studied objects in combinatorics and finite geometry with close connections or equivalences to orthogonal arrays, mutually orthogonal Latin squares, transversal designs, and projective planes as discussed in [6].
@@
A large number of constructions and open problems are associated with $(M+1,S)$-nets and can be found in [6]. Fortunately, for our practically-oriented purposes, linear-algebraic permutations over conveniently chosen rings already provide enough flexibility and are trivial to characterize. In particular, we have the following:
@@
We emphasize that in Examples 1 and 2, we must choose $S$ so that $M\leq\mathsf{lpf}(S)$ otherwise an $(M+1,S)$-net is not formed. If $S=p$ a prime number, then this condition becomes that $M\leq S$. If $M=S$ is required, then Example 3 provides more flexibility in the admissible values of $S$ which can be a prime power.
@@ II-E Inter-Block Scattering via Golomb Rulers
An order $M+1$ ruler $(d_{0},d_{1},\dots,d_{M})$ is a Golomb ruler if all differences between distinct marks $d_{k_{1}}-d_{k_{2}}$ where $k_{1},k_{2}\in[M+1]$ and $k_{1}\neq k_{2}$, are distinct. The length of a Golomb ruler is its largest such difference.
Unless otherwise stated, we will assume that all Golomb rulers are normalized so that $0=d_{0}<d_{1}<\dots<d_{M}$. The length of a Golomb ruler is then simply its largest mark $d_{M}$. The set of distances of a Golomb ruler is the set of its positive differences
@@
which is necessarily of cardinality $\binom{M+1}{2}=\frac{(M+1)M}{2}$ by the distinctness of these differences. This implies that the length $d_{M}$ of an order $M+1$ Golomb ruler satisfies the trivial bound
@@
with equality if and only if its set of distances is precisely $\{1,2,\dots,(M+1)M/2\}$. A Golomb ruler for which the trivial bound holds with equality is said to be perfect.
@@
We provide for the reader’s convenience optimal Golomb rulers of order up to $M+1=10$ in Table I. Observe that the first four optimal Golomb rulers are verified to be optimal by seeing that they are perfect. Beyond these cases, optimal Golomb rulers are typically found by computer search. Optimal Golomb rulers of order up to $M+1=20$ are compiled in [6, Part VI: Ch. 19]. While it is an open problem to exhibit optimal Golomb rulers of every order, Golomb rulers of length $d_{M}<(M+1)^{2}$ have been found for all $M+1<65000$ in [18]. Moreover, an explicit construction of Golomb rulers with cubic length in arbitrary order $M+1$ can also be found in [18].